# An iterative method for the solution of Laplace-like equations in high and very high space dimensions

### v3

@@ 5 The iterative procedure
In this case, the iteration error possesses the radial-angular representation
@@
Since the function $\phi$ takes the constant value $|U|_{2,s}^{2}$ in the given case, this proves the error representation. Convergence is shown as in the proof of Theorem 5.3. ∎
This lemma may not be very interesting on its own, but once again shows that not much is lost in the error estimate (17) and that the prefactors
@@
cannot really be improved. Moreover, it shows that these prefactors tend to zero for optimally or near optimally chosen polynomials $P_{k}$ as $k$ goes to infinity.
@@
This is where our considerations from the previous two sections come into play. We have seen that the values $\|T^{t}\omega\|^{2}$ are approximately normally distributed, with a variance $V$ and a standard deviation $\sigma=\sqrt{V}$ that tend to zero in almost all cases as the dimensions increase. This justifies replacing the actual distribution (9) by the corresponding normal distribution with the density
@@
Table 1: The reduction factors $M_{k}(\sigma)$ assigned to the approximate density (24) for the first $12$ iterations.
@@
The only matrices $T^{t}$ for which the distribution of the values $\|T^{t}\eta\|^{2}$ is explicitly known and available for comparison are the orthogonal projections studied in Sect. 4. The densities of such distributions can be transformed to the weight functions associated with Jacobi polynomials. The orthogonal polynomials assigned to the rescaled variants of these matrices can therefore be expressed in terms of Jacobi polynomials. Details can be found in the appendix. For smaller dimensions, the resulting values (23) tend to zero much faster than the values that one obtains approximating the actual density by the density of a normal distribution. Not surprisingly, the values approach each other as the dimensions increase.
@@
For the matrices $T$ from Sect. 3 assigned to undirected interaction graphs, the angular distribution of the values $\|T^{t}\omega\|^{2}$ is often very similar to that in the case of rescaled orthogonal projections. In fact, it differs in very many cases only very slightly from a to the expected value one rescaled beta distribution. Recall that beta distributions depend on two parameters $\alpha,\beta>0$ and possess the density
$g(t)=\frac{\Gamma(\alpha+\beta)}{\Gamma(\alpha)\Gamma(\beta)}\;t^{\alpha-1}(1-t)^{\beta-1}.$ (28)
The density of their to the expected value one rescaled counterpart is
$f(t)=\frac{\alpha}{\alpha+\beta}\,g\bigg{(}\frac{\alpha\,t}{\alpha+\beta}\bigg{)}$ (29)
on the interval $0<t<(\alpha+\beta)/\alpha$ and zero otherwise.
Lemma 11
The variance $V$ and the third-order central moment $Z$ of the rescaled beta distribution with the density (29) are connected to each other through
${}-\frac{1-V}{2}<\frac{Z}{2V}<V.$ (30)
In terms of $V$ and $Z$, the parameters $\alpha$ and $\beta$ are given by
$\alpha=2\;\frac{Z+V(1-V)}{4V^{2}-(1-V)Z},\quad\beta=\frac{V\alpha(\alpha+1)}{1-V\alpha}.$ (31)
For any given $V>0$ and $Z$ that satisfy the condition (30), conversely there exists a rescaled beta distribution with variance $V$ and third-order central moment $Z$.
Proof
The variance and the third-order central moment of a rescaled beta distribution with the density (29) possess the explicit representation
$V=\frac{\beta}{\alpha(\alpha+\beta+1)},\quad Z=\frac{2\beta(\beta-\alpha)}{\alpha^{2}(\alpha+\beta+1)(\alpha+\beta+2)}.$
From the representation of the variance $V$
$(1-V\alpha)\beta=V\alpha(\alpha+1)$
follows. As $\beta$ and the right-hand side of this identity are positive, $1-V\alpha$ is positive, too, and we obtain the given representation of $\beta$ in terms of $\alpha$ and $V$. If one inserts this into the representation of $Z$, one is led to the identity
$\frac{2V\alpha-(1-V)}{\alpha+2-V\alpha}=\frac{Z}{2V}$
and finally to the representation of $\alpha$. Because the left-hand of this identity grows as function of $\alpha\geq 0$ strictly and $0<\alpha<1/V$ holds, (30) follows. As one can reverse the entire process, the proposition is proved. ∎
Let $T$ be an $(n\times m)$-matrix $T$ with the singular values $\sigma_{1},\ldots,\sigma_{m}$ and let
$A_{1}=\sum_{i=1}^{m}\sigma_{i}^{2},\quad A_{2}=\sum_{i=1}^{m}\sigma_{i}^{4},\quad A_{3}=\sum_{i=1}^{m}\sigma_{i}^{6}$ (32)
be the traces of the matrices $S=T^{t}T$, $S^{2}$, and $S^{3}$. Then the variance and the third-order central moment of the distribution of the values $\|T^{t}\omega\|^{2}$ are given by
$V=\frac{2n\,A_{2}-2A_{1}^{2}}{n^{2}(n+2)},\quad Z=\frac{16A_{1}^{3}-24n\,A_{1}A_{2}+8n^{2}A_{3}}{n^{3}(n+2)(n+4)}.$ (33)
For the matrices $T$ assigned to large classes of graphs, these quantities satisfy the condition (30) in general or with very few exceptions and the resulting densities (29) agree almost perfectly with the actual densities. The values (23) assigned to these densities thus reflect the convergence behavior presumably at least as well as the values resulting from the given approximation by a simple Gauss function. In a sense, this reverses the argumentation and emphasizes the role of the model problem analyzed in Sect. 4. From this point of view, the beta distributions are at the beginning and the normal distributions only the limit case for higher dimensions.
@@
To check with which relative error the function (35) approximates $1/r$ on a given interval $1\leq r\leq R$, thus one has to check how well the function $\phi$ approximates the constant $1$ on the interval $0\leq s\leq\ln R$.
@@
$\sim 4\pi h^{-1/2}\mathrm{e}^{-\pi^{2}/h}$ (38)
as $h$ goes to zero. High relative accuracy on large intervals $1\leq r\leq R$ is a considerably stronger requirement than high absolute accuracy. The approximation of $1/r$ with minimum absolute error has been studied in Braess-Hackbusch , Braess-Hackbusch_2 , and Hackbusch_2 . The technique in Hackbusch_2 can be used to compute the approximations with the least relative error. On the given interval, one gains a factor of just over five with a sum of $50$ exponential functions. Hackbusch has compiled the best approximations for a large number of intervals and sums of exponential functions of very different length. The data can be downloaded from the website of the Max Planck Institute in Leipzig Hackbusch_3 .
Figure 4: The rescaled function $s\to\phi(s\ln 10)$ approximating $1$ for $h=1$, $k_{1}=-2$, and $k_{2}=50$.
@@
The goal in the background is to apply the developed techniques to approximate inverse iterations for the calculation of the ground state of the electronic Schrödinger equation; see Scholz-Yserentant for more details. Electronic wave functions are subject to the Pauli principle. Taking the spin into account, they are antisymmetric with respect to the exchange of the electrons, and if one considers the different spin components of the wave functions separately, they are antisymmetric with respect to the exchange of the electrons with the same spin. The following considerations show that such symmetry properties are preserved in the course of the calculations.

### v5

@@ 4 The iterative procedure
In radial-angular representation, the iteration error is given by
@@
Convergence follows as in the proof of Theorem 4.3. If $U$ is rotationally symmetric, the function $\phi$ takes the constant value $|U|_{2,\,s}^{2}$ and (20) is proven. ∎
Because equality holds in (20), this once again shows that little is lost in the error estimate (19) and that the prefactors
@@
cannot really be improved and are de facto optimal.
@@
In the previous section, we have seen that the values $\|T^{t}\eta\|^{2}$ are approximately normally distributed, with a variance $V$ and a standard deviation $\sigma=\sqrt{V}$ that tend to zero in almost all cases as the dimensions increase. This justifies replacing the actual distribution (2) by the corresponding normal distribution with the density
@@
Table 1: The reduction factors $M_{k}(\sigma)$ assigned to the approximate density (25) for the first $12$ iterations
The only matrices $T^{t}$ for which the distribution density (9) is explicitly known and available for comparison are the rescaled versions of the orthogonal projections from $\mathbb{R}^{n}$ to $\mathbb{R}^{m}$, $(m\times n)$-matrices with one as the only singular value. By Theorem 3.1, the densities (9) assigned to such orthogonal projections take the value
$g(t)=\frac{1}{B(\alpha,\beta)}\;t^{\alpha-1}(1-t)^{\beta-1}$ (29)
on the interval $10<t<$ 1 and vanish outside of it, where $\alpha$ and $\beta$ in this case are given by (10). The to the expected value one rescaled counterpart of such densities is
$f(t)=\frac{\alpha}{\alpha+\beta}\,g\bigg{(}\frac{\alpha\,t}{\alpha+\beta}\bigg{)}.$ (30)
@@
These densities can be transformed to the weight functions associated with Jacobi polynomials. The orthogonal polynomials assigned to them can thus be expressed in terms of Jacobi polynomials. The details can be found in the appendix. For smaller dimensions, the resulting values (24) tend to zero much faster than the values that one gets approximating the actual densities by the densities of the corresponding normal distributions. This is likely due the behavior of the given beta distributions for small $\delta$, a behavior that is common to all distributions (2) by Theorem 3.1. However, the values rapidly approach each other as the dimensions increase.
@@
For the matrices $T$ from Sect. 3 assigned to undirected interaction graphs, the angular distribution of the values $\|T^{t}\omega\|^{2}$ is often very similar to that in the case of rescaled orthogonal projections. In fact, it differs in many cases only very slightly from a to the expected value one rescaled beta distribution. The variance of a rescaled beta distribution with the density (30) is
$V=\frac{\beta}{\alpha(\alpha+\beta+1)}.$ (31)
The parameters $\alpha$ and $\beta$ and the variance $V$ are connected via the relation
$(1-V\alpha)\beta=V\alpha(\alpha+1).$ (32)
Thus, such a beta distribution with given parameter $\alpha$ and given variance $V$ exists if and only if $01-V\alpha>$ 0. In this case, it is uniquely determined and $\beta$ is given by
$\beta=\frac{V\alpha(\alpha+1)}{1-V\alpha}.$ (33)
If one sets $\alpha$ to the value $m/2$ given by Theorem 3.1, the compatibility condition
$V<\frac{2}{m}$ (34)
is satisfied for the matrices $T$ assigned to large classes of graphs, without exception or with high probability. This follows from the representation (26) of the variance in terms of the vertex degrees. Examples are the matrices assigned to regular graphs, whose vertices all have the same number of neighbors, or the matrices assigned to random graphs with a fixed number of vertices and edges. In such cases, the resulting densities (30) are almost identical to the actual densities. The values (24) assigned to them therefore reflect the convergence behavior presumably at least as well as the values resulting from the given approximation by a simple Gauss function. Figure 3 compares the distribution density originating from the matrix assigned to a small random graph with $m=32$ vertices and $n-m=96$ edges with its approximation by the density of the corresponding normal distribution and that by the density of the given rescaled beta distribution. The variance is $V=941/16640$ in this example. The difference is striking, but soon disappears for larger graphs.
Figure 3: Comparison of the approximation by a rescaled beta distribution and a normal distribution
@@
To determine with which relative error the function (36) approximates $1/r$ on a given interval $1\leq r\leq R$, thus one must check how well the function $\phi$ approximates the constant $1$ on the interval $0\leq s\leq\ln R$.
@@
$\sim 4\pi h^{-1/2}\mathrm{e}^{-\pi^{2}/h}.$ (39)
High relative accuracy on large intervals $1\leq r\leq R$ is a considerably stronger requirement than high absolute accuracy. The approximation of $1/r$ with minimum absolute error has been studied in Braess-Hackbusch , Braess-Hackbusch_2 , and Hackbusch_2 . The technique in Hackbusch_2 can be used to compute the approximations with the least relative error. On the given interval, one gains a factor of just over five with a sum of $50$ exponential functions. Hackbusch has compiled the best approximations for a large number of intervals and sums of exponential functions of very different length. The data can be downloaded from the website of the Max Planck Institute in Leipzig Hackbusch_3 .
Figure 4: The rescaled function $s\to\phi(s\ln 10)$ approximating $1$ for $h=1$, $k_{1}=-2$, and $k_{2}=50$
@@
Remark
This is a largely rewritten version of the paper [Numer. Math. (2024) 156:777–811] of the same title. A central new result is Theorem 3.1. The present version uses two new scales of norms that directly measure the smoothness of the trace functions and fit better into the given framework.
@@
The goal in the background is the application of the here developed techniques to approximate inverse iterations for the calculation of the ground state of the electronic Schrödinger equation; see Scholz-Yserentant for more details. Electronic wave functions are subject to the Pauli principle. Taking the spin into account, they are antisymmetric with respect to the exchange of the electrons, and if one considers the different spin components of the wave functions separately, they are antisymmetric with respect to the exchange of the electrons with the same spin. The following considerations show that such symmetry properties are preserved in the course of the calculations.